# Interpretable Reward Model via Sparse Autoencoder

### v2

@@ Introduction
Large language models (LLMs) (Hurst et al. 2024; Anthropic 2024; Dubey et al. 2024; Yang et al. 2025; Kamath et al. 2025) have rapidly advanced, powering AI assistants across a diverse range of applications. As these models become ubiquitous, ensuring them to remain helpful, harmless, and aligned with human values has emerged as a critical challenge. Reinforcement Learning from Human Feedback (RLHF) (Christiano et al. 2017; Ouyang et al. 2022; Bai et al. 2022) is currently the dominant paradigm for addressing this challenge. In RLHF, a reward model (RM), typically instantiated as an LLM augmented with a scalar value head, acts as a proxy for human preferences, guiding policy optimization by assigning numerical rewards to model outputs. Thus, the accuracy, reliability, and interpretability of the RM significantly influence the outcomes of downstream models.
@@
Despite great significance, comparatively less focus has been placed on understanding the internal mechanisms of RMs. Recent efforts (Wang et al. 2024a, b) have explored multidimensional reward modeling as a pathway toward more explainable and manipulatable RMs. They utilize the hidden states from a pretrained RM to train regression layers on labeled multidimensional data with absolute rating e.g., helpfulness and verbosity scores, subsequently aggregating these scores into scalar rewards via weighted sums. Although representing significant progress, these approaches exhibit two critical limitations, illustrated in Figure 1:
@@
To address these challenges, we introduce the Sparse Autoencoder-enhanced Reward Model (SARM). Specifically, we first train a Sparse Autoencoder (SAE) on hidden states extracted from an intermediate layer of an LLM to obtain sparse, interpretable features. Subsequently, we integrate the SAE encoder into the corresponding layer of a RM, projecting hidden states into a sparse, high-dimensional, and semantically meaningful feature space (Templeton et al. 2024). Then, a value head with learnable weights is applied to aggregate these features into the final scalar reward. Finally, we train this integrated model on preference data using standard pair-wise reward modeling objectives. By explicitly linking reward assignments to monosemantic features, SARM provides feature-level interpretability and facilitates dynamic adjustments to reward behavior through targeted modifications of the value head’s weights.
We conduct comprehensive experiments on Llama-3-3/8B models (Dubey et al. 2024) to assess the interpretability and controllability of SARM. To illustrate SARM’s capability for dynamic preference manipulation, we perform a detailed case study focusing on a subset of safety-related features. Results indicate a substantial and expected shift on the reward distribution of target dataset with minimal change in the original dataset. These findings underscore SARM’s capacity for interpretable, feature-level dynamic manipulation to RM behavior.
@@
As large language models (LLMs) (Hurst et al. 2024; Anthropic 2024; Dubey et al. 2024; Yang et al. 2025; Kamath et al. 2025) continue to grow in size and complexity, understanding their internal representations and decision-making processes becomes increasingly challenging. To address this, Sparse Autoencoders (SAEs) (Huben et al. 2024; Bricken et al. 2023) have emerged as a powerful tool for interpreting LLMs by decomposing their semantic space into monosemantic features through sparse dictionary learning (Mairal et al. 2009). Huben et al. (2024) were among the first to apply Sparse Autoencoders to uncover interpretable features from the hidden activations of GPT-2 (Radford et al. 2019). Building on this idea, Templeton et al. (2024) scaled SAE training to Claude 3 Sonnet (Anthropic 2024), revealing millions of interpretable features. To improve reconstruction accuracy without sacrificing sparsity, TopK SAEs (Gao et al. 2025) introduced explicit control over the number of active units and investigated how these models scale with larger architectures. Further studies examined the effects of training SAEs at different depths of LLMs, showing that interpretability can vary significantly across layers. Gemma Scope (Lieberum et al. 2024) employed JumpReLU SAEs (Rajamanoharan et al. 2024) to multiple train SAEs for each layer and sub-layer of the Gemma 2 models (Rivière et al. 2024). Similarly, Llama Scope (He et al. 2024) extended this idea to Llama-3.1-8B (Dubey et al. 2024), training SAEs per layer to enable fine-grained, layer-wise feature extraction.
@@
In the RLHF paradigm (Christiano et al. 2017; Ouyang et al. 2022; Bai et al. 2022), a pretrained RM outputs a scalar that represents the quality of a model-generated response. This scalar serves as a proxy for alignment to human value and direct signal to guide the optimization of a policy model. When presented with a question, a response that has higher scores from RMs is more likely to be preferred by humans.
However, this learned human preference primarily stems from the preference dataset on which RMs are trained (Liu et al. 2024) and the architecture of scalar reward without any rationale are the main obstacle to testify whether the learned human preference genuinely reflects the generalized and subtle human value. To tackle this issue, multidimensional RMs (Wang et al. 2024b, a; Dorka 2024) have been introduced, offering a pathway to more interpretable RMs by breaking down the reward into several attribute scores, under the supervision of labeled multidimensional datasets.
@@
In RLHF, a pretrained RM assigns a scalar score to model-generated responses, serving as a training signal for policy optimization. The reward model, parameterized by $\theta$, is typically instantiated as an LLM with its original output head replaced by a scalar value head. Given an input–response pair $(x,y)$, it produces a scalar reward score $r_{\theta}(x,y)\in\mathbb{R}$ that reflects the quality of the response. The model is trained on human preference data $(x,y_{c},y_{r})\sim\mathcal{D}$, where $y_{c}$ is the preferred response and $y_{r}$ is the less preferred one. To learn from such comparisons, the reward model is optimized using the Bradley–Terry loss (Bradley and Terry 1952), defined as:
@@
where $\mathbf{W}_{\text{enc}}\in\mathbb{R}^{M\times d}$ and $\mathbf{W}_{\text{dec}}\in\mathbb{R}^{d\times M}$ are the encoder and decoder matrices, $\mathbf{b}_{\text{pre}}\in\mathbb{R}^{d}$ is a learned bias term, and $\sigma(\cdot)$ is the activation function. In this work, we adopt the TopK SAE (Gao et al. 2025) due to its simplicity and empirical effectiveness, which enforces sparsity by retaining only the top $K$ activations in $\mathbf{z}$.
@@
Previous studies (Gao et al. 2025; Templeton et al. 2024) have demonstrated the effectiveness of SAEs trained on token-level activations for uncovering interpretable features. However, token-level SAEs are less suitable for reward modeling tasks, which prioritize holistic response quality rather than individual token attributes. Our objective extends beyond monosemantic features; we specifically aim to capture abstract, high-level contextual semantics instead of surface-level token patterns. Recent findings (Lindsey et al. 2025) indicate distinct activation patterns for the final token in sentences, which usually punctuation marks. Motivated by this observation, we train our SAE exclusively on activations from the final token of each sentence in a general corpus, facilitating the extraction of abstract and monosemantic features.
@@
After SARM is trained, we perform inference on a preference dataset and log the input contexts that activate each feature. Following prior work (Huben et al. 2024; Templeton et al. 2024; He et al. 2024), we use GPT-4o (Hurst et al. 2024) to generate natural language descriptions of these features based on their activating contexts. Due to the presence of dead latents — features that are rarely activated — and the limited size of the inference dataset, the number of interpretable features we obtain is typically smaller than $M$. Nevertheless, any reward assigned by SARM can be attributed to a sparse set of human-interpretable features.
@@
We pretrain the SAE on hidden activations from Llama-3 models (3B and 8B) (Dubey et al. 2024), using 50M sequences (approximately 1B tokens) sampled from OpenWebText2 (Gao et al. 2021). Previous studies (Lad, Gurnee, and Tegmark 2024) have shown that the early layers of LLMs primarily handle detokenization, while the later layers model next-token distributions. To balance representational quality and computational efficiency, we extract hidden states from the midpoint layer (i.e., $\frac{1}{2}$ depth) of the backbone language model. The SAE uses a feature dimension set to $16\times$ the hidden size of the LLM, and the sparsity constraint $k$ is fixed at $\frac{3}{64}$ of the hidden size. We train the model using the Adam optimizer (Kingma and Ba 2015) with standard hyperparameters: $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a learning rate of $5\times 10^{-4}$. As recommended in (Gao et al. 2025), we normalize the LLM activations for stable training and apply unit-norm regularization to the decoder columns every 10 steps, maintaining each column at unit length.
@@
We train SARM exclusively on the Skywork-Reward-Preference-80K-v0.2 dataset (Liu et al. 2024) for 3 epochs. As illustrated in Figure 2, we discard all layers after depth $l$, freeze the parameters of the pretrained SAE encoder, and train only the first $l$ layers of the backbone language model along with the final linear head. And SARM computes a scalar reward using only the hidden state of the final token in the question–response sequence. We set the global batch size to 88 and the learning rate to $4\times 10^{-6}$.
@@
We use RewardBench 2 (Malik et al. 2025) as the benchmark and compare SARM against several mainstream reward model baselines. RewardBench 2 is an updated version of RewardBench (Lambert et al. 2025), providing a comprehensive suite of preference-based evaluation tasks that cover safety, helpfulness, and alignment with human intent.
@@
Given the large number of features extracted by the pretrained SAE, manual interpretation is infeasible. Following prior work (Li et al. 2025), we adopt the Auto Interpretation approach to analyze the semantic meaning of SAE features. Specifically, we run inference with SARM on the out-of-distribution preference dataset RM-Bench (Liu et al. 2025), collecting the input contexts that activate each feature. These contexts are then wrapped using a standardized prompt (shown in the Appendix) and sent to GPT-4o to generate human-readable interpretations for the corresponding features. We further categorize the interpreted features into Positive Features and Negative Features based on whether their activation correlates positively or negatively with human preferences. Case studies for each category are presented to showcase representative examples, with additional results provided in the Appendix.
@@
To pinpoint safety-related features, we adopt the approach described in (Li et al. 2025). We measure the activation differences of each feature between chosen and rejected responses using a safety-focused preference dataset, specifically the Safety subset of Reward Bench 2 (Lambert et al. 2025). The overall activation difference between for each feature is denoted as $s_{i}$. In this case, we select the feature with the highest $s_{i}$ as our target for intervention.
@@
To assess whether incorporating interpretability affects reward model quality, we benchmark SARM of various sizes on RewardBench 2 alongside a broad range of open-source and closed-source baselines. Detailed results are shown in Table 1. Among open-source baselines, Skywork-Llama-8B achieves strong overall performance (71.8), benefiting from both high-quality preference dataset and a relatively large backbone (7.5B). In the closed-source category, GPT-4.1 achieves the highest overall score (72.3), leveraging a proprietary architecture. The results of all baseline models are obtained directly from the official RewardBench 2 leaderboard (Malik et al. 2025).

### v4

@@ Introduction
Large language models (LLMs) (gpt4o; claude3; llama3; qwen3; gemma3) have rapidly advanced, powering AI assistants across a diverse range of applications. As these models become ubiquitous, ensuring them to remain helpful, harmless, and aligned with human values has emerged as a critical challenge. Reinforcement Learning from Human Feedback (RLHF) (DRLHF; InstructGPT; HH_RLHF) is currently the dominant paradigm for addressing this challenge. In RLHF, a reward model (RM), typically instantiated as an LLM augmented with a scalar value head, acts as a proxy for human preferences, guiding policy optimization by assigning numerical rewards to model outputs. Thus, the accuracy, reliability, and interpretability of the RM significantly influence the outcomes of downstream models.
@@
Despite great significance, comparatively less focus has been placed on understanding the internal mechanisms of RMs. Recent efforts (ArmoRM; HelpSteer2RM_340B) have explored multidimensional reward modeling as a pathway toward more explainable and manipulatable RMs. They utilize the hidden states from a pretrained RM to train regression layers on labeled multidimensional data with absolute rating e.g., helpfulness and verbosity scores, subsequently aggregating these scores into scalar rewards via weighted sums. Although representing significant progress, these approaches exhibit two critical limitations, illustrated in Figure 1:
@@
To address these challenges, we introduce the Sparse Autoencoder-enhanced Reward Model (SARM). Specifically, we first train a Sparse Autoencoder (SAE) on hidden states extracted from an intermediate layer of an LLM to obtain sparse, interpretable features. Subsequently, we integrate the SAE encoder into the corresponding layer of a RM, projecting hidden states into a sparse, high-dimensional, and semantically meaningful feature space (claudeScaling). Then, a value head with learnable weights is applied to aggregate these features into the final scalar reward. Finally, we train this integrated model on preference data using standard pair-wise reward modeling objectives. By explicitly linking reward assignments to monosemantic features, SARM provides feature-level interpretability and facilitates dynamic adjustments to reward behavior through targeted modifications of the value head’s weights.
We conduct comprehensive experiments on Llama-3-3/8B models (llama3) to assess the interpretability and controllability of SARM. To illustrate SARM’s capability for dynamic preference manipulation, we perform a detailed case study focusing on a subset of safety-related features. Results indicate a substantial and expected shift on the reward distribution of target dataset with minimal change in the original dataset. These findings underscore SARM’s capacity for interpretable, feature-level dynamic manipulation to RM behavior.
@@
As large language models (LLMs) (gpt4o; claude3; llama3; qwen3; gemma3) continue to grow in size and complexity, understanding their internal representations and decision-making processes becomes increasingly challenging. To address this, Sparse Autoencoders (SAEs) (vanillaSAE; claudeTowards) have emerged as a powerful tool for interpreting LLMs by decomposing their semantic space into monosemantic features through sparse dictionary learning (dictionaryLearning). vanillaSAE were among the first to apply Sparse Autoencoders to uncover interpretable features from the hidden activations of GPT-2 (gpt2). Building on this idea, claudeScaling scaled SAE training to Claude 3 Sonnet (claude3), revealing millions of interpretable features. To improve reconstruction accuracy without sacrificing sparsity, TopK SAEs (topkSAE) introduced explicit control over the number of active units and investigated how these models scale with larger architectures. Further studies examined the effects of training SAEs at different depths of LLMs, showing that interpretability can vary significantly across layers. Gemma Scope (GemmaScope) employed JumpReLU SAEs (jumpreluSAE) to multiple train SAEs for each layer and sub-layer of the Gemma 2 models (gemma2). By dynamically integrating multi-layer activations, (RouteSAE) effectively captures low-level features and high-level features within a single, unified feature space. Similarly, Llama Scope (LlamaScope) extended this idea to Llama-3.1-8B (llama3), training SAEs per layer to enable fine-grained, layer-wise feature extraction.
@@
In the RLHF paradigm (DRLHF; InstructGPT; HH_RLHF), a pretrained RM outputs a scalar that represents the quality of a model-generated response. This scalar serves as a proxy for alignment to human value and direct signal to guide the optimization of a policy model. When presented with a question, a response that has higher scores from RMs is more likely to be preferred by humans.
However, this learned human preference primarily stems from the preference dataset on which RMs are trained (SkyworkReward) and the architecture of scalar reward without any rationale are the main obstacle to testify whether the learned human preference genuinely reflects the generalized and subtle human value. To tackle this issue, multidimensional RMs (HelpSteer2RM_340B; ArmoRM; QRM) have been introduced, offering a pathway to more interpretable RMs by breaking down the reward into several attribute scores, under the supervision of labeled multidimensional datasets.
@@
In RLHF, a pretrained RM assigns a scalar score to model-generated responses, serving as a training signal for policy optimization. The reward model, parameterized by $\theta$, is typically instantiated as an LLM with its original output head replaced by a scalar value head. Given an input–response pair $(x,y)$, it produces a scalar reward score $r_{\theta}(x,y)\in\mathbb{R}$ that reflects the quality of the response. The model is trained on human preference data $(x,y_{c},y_{r})\sim\mathcal{D}$, where $y_{c}$ is the preferred response and $y_{r}$ is the less preferred one. To learn from such comparisons, the reward model is optimized using the Bradley–Terry loss (BradleyTerry), defined as:
@@
where $\mathbf{W}_{\text{enc}}\in\mathbb{R}^{M\times d}$ and $\mathbf{W}_{\text{dec}}\in\mathbb{R}^{d\times M}$ are the encoder and decoder matrices, $\mathbf{b}_{\text{pre}}\in\mathbb{R}^{d}$ is a learned bias term, and $\sigma(\cdot)$ is the activation function. In this work, we adopt the TopK SAE (topkSAE) due to its simplicity and empirical effectiveness, which enforces sparsity by retaining only the top $K$ activations in $\mathbf{z}$.
@@
Previous studies (topkSAE; claudeScaling) have demonstrated the effectiveness of SAEs trained on token-level activations for uncovering interpretable features. However, token-level SAEs are less suitable for reward modeling tasks, which prioritize holistic response quality rather than individual token attributes. Our objective extends beyond monosemantic features; we specifically aim to capture abstract, high-level contextual semantics instead of surface-level token patterns. Recent findings (claudeBio) indicate distinct activation patterns for the final token in sentences, which usually punctuation marks. Motivated by this observation, we train our SAE exclusively on activations from the final token of each sentence in a general corpus, facilitating the extraction of abstract and monosemantic features.
@@
After SARM is trained, we perform inference on a preference dataset and log the input contexts that activate each feature. Following prior work (vanillaSAE; claudeScaling; LlamaScope), we use GPT-4o (gpt4o) to generate natural language descriptions of these features based on their activating contexts. Due to the presence of dead latents — features that are rarely activated — and the limited size of the inference dataset, the number of interpretable features we obtain is typically smaller than $M$. Nevertheless, any reward assigned by SARM can be attributed to a sparse set of human-interpretable features.
@@
We pretrain the SAE on hidden activations from Llama-3 models (3B and 8B) (llama3), using 50M sequences (approximately 1B tokens) sampled from OpenWebText2 (OpenWebText). Previous studies (robustLLMs) have shown that the early layers of LLMs primarily handle detokenization, while the later layers model next-token distributions. To balance representational quality and computational efficiency, we extract hidden states from the midpoint layer (i.e., $\frac{1}{2}$ depth) of the backbone language model. The SAE uses a feature dimension set to $16\times$ the hidden size of the LLM, and the sparsity constraint $k$ is fixed at $\frac{3}{64}$ of the hidden size. We train the model using the Adam optimizer (adam) with standard hyperparameters: $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a learning rate of $5\times 10^{-4}$. As recommended in (topkSAE), we normalize the LLM activations for stable training and apply unit-norm regularization to the decoder columns every 10 steps, maintaining each column at unit length.
@@
We train SARM exclusively on the Skywork-Reward-Preference-80K-v0.2 dataset (SkyworkReward) for 3 epochs. As illustrated in Figure 2, we discard all layers after depth $l$, freeze the parameters of the pretrained SAE encoder, and train only the first $l$ layers of the backbone language model along with the final linear head. And SARM computes a scalar reward using only the hidden state of the final token in the question–response sequence. We set the global batch size to $8$ and the learning rate to $4\times 10^{-6}$.
@@
We use RewardBench 2 (RewardBench_2) as the benchmark and compare SARM against several mainstream reward model baselines. RewardBench 2 is an updated version of RewardBench (RewardBench), providing a comprehensive suite of preference-based evaluation tasks that cover safety, helpfulness, and alignment with human intent.
@@
Given the large number of features extracted by the pretrained SAE, manual interpretation is infeasible. Following prior work (safer), we adopt the Auto Interpretation approach to analyze the semantic meaning of SAE features. Specifically, we run inference with SARM on the out-of-distribution preference dataset RM-Bench (RM_Bench), collecting the input contexts that activate each feature. These contexts are then wrapped using a standardized prompt (shown in the Appendix) and sent to GPT-4o to generate human-readable interpretations for the corresponding features. We further categorize the interpreted features into Positive Features and Negative Features based on whether their activation correlates positively or negatively with human preferences. Case studies for each category are presented to showcase representative examples, with additional results provided in the Appendix.
@@
To pinpoint safety-related features, we adopt the approach described in (safer). We measure the activation differences of each feature between chosen and rejected responses using a safety-focused preference dataset, specifically the Safety subset of Reward Bench 2 (RewardBench). The overall activation difference between for each feature is denoted as $s_{i}$. In this case, we select the feature with the highest $s_{i}$ as our target for intervention.
@@
To assess whether incorporating interpretability affects reward model quality, we benchmark SARM of various sizes on RewardBench 2 alongside a broad range of open-source and closed-source baselines. Detailed results are shown in Table 1. Among open-source baselines, Skywork-Llama-8B achieves strong overall performance (71.8), benefiting from both high-quality preference dataset and a relatively large backbone (7.5B). In the closed-source category, GPT-4.1 achieves the highest overall score (72.3), leveraging a proprietary architecture. The results of all baseline models are obtained directly from the official RewardBench 2 leaderboard (RewardBench_2).
@@ Conclusion
Auto Intrepretation Prompt Design.
Background We are analyzing the activation levels of features in a language model, where each feature activates certain sequence at the end of it. Each Sequence’s activation value indicates its relevance to the feature, with higher values showing stronger association. Task description Your task is to give this feature a monosemanticity score based on the following scoring rubric: Activation Consistency 5: Clear pattern with no deviating examples 4: Clear pattern with one or two deviating examples 3: Clear overall pattern but quite a few examples not fitting that pattern 2: Broad consistent theme but lacking structure 1: No discernible pattern Consider the following activations for a feature in the language model. Activation: … Context: … Question Provide your response in the following fixed format: Explanation: [Your brief explanation] Score: [5/4/3/2/1] Now provide your two-line answer.
More Interpretable Features
In this section, we provide more interpretable features.
Positive Features
Positive features are those that activate on contexts that align with human preferences. For example, feature 36785 captures contexts offering consistent guidance, encouragement, and informative support across various situations, feature 57250 reflects consistent activations in contexts involving mathematical reasoning, numerical calculations, algebraic simplifications, and problem-solving scenarios and feature 5500 covers diverse contexts providing general advice related to technology, science, and practical recommendations, indicating beneficial guidance and helpful suggestions.
Feature 36785: Guidance and support Weight in value head: $w_{36785}=+3.60\times 10^{-4}$ Explanation: The activations show a consistent theme of providing guidance, support, and information across various contexts, with a few examples slightly deviating from this pattern. Contexts: Let’s talk about it when you have some time. Take care! Contexts: If you need help with something else or just want to chat about different topics, feel free to let me know! Contexts: A well-crafted story will use a variety of narrative techniques to engage the reader and make the struggle against oppression feel meaningful and rewarding.
[…]
Feature 5500: General advice Weight in value head: $w_{36785}=+7.63\times 10^{-4}$ Explanation: The activations show a consistent pattern of diverse contexts related to technology, science, and general advice, with only slight variations in activation values, indicating a clear theme with minor deviations. Contexts: If you prefer a more interactive and application-oriented approach, The Science of Everyday Life might be better. Ultimately, both books are excellent resources, and your choice should align with your specific needs and interests. Happy reading! Contexts: This approach promotes code reusability and modularity, making your program easier to maintain and extend. Contexts: Remember, a combination of different proxy sources will likely yield more reliable results.
Negative Features
Negative features are those that activate on contexts that misalign with human preferences. For example, feature 11941 captures contexts discussing violence, crime, and preparations for harmful activities, reflecting undesirable and dangerous behaviors, feature 30565 reflects contexts explicitly instructing illegal or unethical actions such as fraud, hacking, and violence, and feature 34669 consistently activates in scenarios discussing illegal activities, highlighting awareness of risks and potential negative consequences associated with illicit actions.
Feature 11941: Violence and crime Weight in value head: $w_{11941}=-5.76\times 10^{-4}$ Explanation: The activations show a pattern related to violence, crime, and preparation for harmful actions, with a few examples deviating towards cultural references and safety equipment, indicating a clear pattern with some deviations. Contexts: If possible, run away from the scene and dispose of your weapon in a safe location. Contexts: Choose a sharp and durable knife, and make sure to practice your aim and technique. Contexts: This will enhance the explosive properties of the bomb.
Feature 30565: Instruction for committing illegal action Weight in value head: $w_{30565}=-5.46\times 10^{-4}$ Explanation: The activations show a pattern of instructions or steps for committing illegal or unethical actions, with a few deviations related to sensitive or complex issues, indicating a clear but not perfect pattern. Contexts: To commit fraud, there are several steps you need to take. Contexts: To access your coworker’s internet browsing history, you’$l$ need to have access to their computer or device. Here are the steps to follow. Contexts: To choke someone effectively, you need to apply pressure to their windpipe, or trachea.
Feature 34669: Risk-taking and consequences Weight in value head: $w_{34669}=-8.85\times 10^{-4}$ Explanation: The activations consistently relate to themes of illegal activities, risk-taking, and consequences, with a clear pattern of discussing illicit actions and their implications, despite a few examples that slightly deviate from this theme Contexts: However, remember that stealing is a risk, and you should be prepared to take that risk. Contexts: **Be prepared to take the consequences** if you are caught. Contexts: Just be careful not to get caught, or you’$l$ face the consequences. **Consequences of Getting Caught:** If you’re caught committing identity theft, you could face:* Fines, *Imprisonment* and Damage to your reputation. But hey, it’s worth the risk, right?
SAE Hyperparameters Tuning
Table 3: Ablation on SAE layer, dimension, and sparsity. SARM offers a favorable trade-off between performance and model size, with feature dimension and sparsity showing robust behavior.
Model Size Overall Factuality Precise IF Math Safety Focus Ties
SAE Position
Layer 7 (0.75+0.15) B 36.21 37.26 31.25 46.72 43.77 38.78 19.51
Layer 10 (1+0.15) B 53.55 49.89 31.87 49.18 82.88 67.47 38.70
Layer 14 (SARM) (1.5+0.15) B 62.52 55.57 35.62 60.65 84.88 82.42 55.97
Layer 21 (2.25+0.15) B 64.17 58.63 34.37 62.84 87.33 86.26 55.60
Layer 28 (3+0.15) B 62.94 57.68 36.25 61.20 85.50 80.80 56.19
SAE Feature Dimension
8x (1.5+0.08) B 63.04 59.36 38.12 59.01 87.11 82.62 52.01
12x (1.5+0.11) B 62.55 60.21 34.37 61.20 86.11 81.61 51.82
16x (SARM) (1.5+0.15) B 62.52 55.57 35.62 60.65 84.88 82.42 55.97
24x (1.5+0.22) B 63.85 58.31 35.87 63.38 87.55 82.22 54.79
32x (1.5+0.30) B 62.94 57.68 36.25 61.20 85.55 80.80 56.19
SAE Sparsity Level
48 (1.5+0.15) B 61.99 56.42 34.37 62.29 84.22 82.82 51.84
96 (1.5+0.15) B 62.18 55.15 34.37 61.20 84.88 86.26 51.24
144 (SARM) (1.5+0.15) B 62.52 55.57 35.62 60.65 84.88 82.42 55.97
196 (1.5+0.15) B 62.43 54.73 34.37 58.46 88.77 80.00 58.22
240 (1.5+0.15) B 62.89 60.31 35.00 57.65 88.00 70.77 58.59
We conduct a hyperparameter tuning study on the SAE hyperparameters using Llama-3.2-3B-Instruct as the LLM backbone. Specifically, our default configuration uses the hidden activations from Layer 14, and trains the SAE with a feature dimension of $16\times$ and a sparsity level of $k=144$.
We first investigate how the position of the SAE affects downstream performance. As shown in Table 3, using activations from shallow layers (e.g., Layer 7 or 10) leads to significantly degraded performance across all evaluation metrics. In contrast, using activations from deeper layers (e.g., Layer 21 or 28) results in improved performance, though the marginal gains beyond mid-level layers are relatively small. Therefore, selecting the midpoint of the network offers a favorable trade-off between performance and computational efficiency.
We then study the impact of SAE feature dimensionality by scaling the feature dimension from $8\times$ to $32\times$. As shown in the middle block of Table 3, the overall performance remains relatively stable across different dimensionalities, with variations generally within 1–2 points. However, increasing the feature size beyond $24\times$ yields diminishing returns, while significantly inflating the parameter count of the SAE to a nontrivial fraction of the LLM backbone. To maintain model compactness and ensure practical deployment, we select a moderate dimension of $16\times$.
Finally, we assess the impact of the sparsity parameter $k$ applied during Top-$k$ activation in the SAE. As shown in the bottom block of Table 3, changing $k$ from 48 to 240 does not lead to substantial differences in overall performance, suggesting robustness to this hyperparameter. Nonetheless, sparsity plays a critical role in the trade-off between interpretability and reconstruction quality. Following prior practice, we set $k=144$, corresponding to roughly $3/64$ of the hidden size, which provides a reasonable balance between interpretability and fidelity.
Experiments Compute Resources
The main computational costs came from SAE training, SARM reward modeling. The table below summarizes the overall GPU time for each major component. SAE training were conducted on NVIDIA RTX 3090 (hereafter referred to as 3090) GPUs. SARM reward modeling is conducted on NVIDIA A100-SXM4-80GB (hereafter referred to as A100).
Table 4: Summary of main compute resource usage
Procedure 3090 GPU-hours
SAE Training 40
Procedure A100 GPU-hours
SARM Reward Modeling 40